OpenStudy (alexacelio1238):

0.000000057 in scientific notation would be

OpenStudy (3mar):

Welcome to Openstudy!

OpenStudy (rcp031):

scientific notation? means is a special way of writing numbers: How to Do it. When the number is smaller than 1, the decimal point has to move to the right, so the power of 10 is negative. Example: 0.0055 is written 5.5 × 10-3 Because 0.0055 = 5.5 × 0.001 = 5.5 × 10-3

KjSaif (kjsaif):

how many zeros are on the right

KjSaif (kjsaif):

@alexacelio1238

KjSaif (kjsaif):

7

KjSaif (kjsaif):

5.7x10^-8

likeabossssssss (likeabossssssss):

welcome to os and plz close this :)